A multiple choice test consists of 8 questions with 

3 answers of each question. A student answers 

each question by rolling a dice and checking 

the first answer if he gets 1 or 2 then second answer 

3 or 4 and third if he gets 5 or 6.to get a distinction 

He/she must secure at least 75% correct answer. 

If there is no negative marking then 

Solve the probability that student secures distinction.



1
Expert's answer
2021-12-27T08:18:53-0500

Solution:

"p=\dfrac13, q=\dfrac23"

The probability of securing a distinction (i.e., getting correct answer of at least 6 of the 8 questions) "=P(X\ge6)=P(X=6)+P(X=7)+P(X=8)"

"=\ ^8C_6(\dfrac13)^6(\dfrac23)^2+\ ^8C_7(\dfrac13)^7(\dfrac23)^1+\ ^8C_8(\dfrac13)^8(\dfrac23)^0
\\=(\dfrac13)^6[\ ^8C_6(\dfrac23)^2+\ ^8C_7(\dfrac13)(\dfrac23)^1+\ ^8C_8(\dfrac13)^2(\dfrac23)^0]
\\=\dfrac{1}{729}[28\times \dfrac{4}{9}+8\times \dfrac{2}{9}+\dfrac{1}{9}]
\\=0.0196"